# 3cm3cm2.3cm2.3cm

### v1

@@ 5.1. Settings
${\mathbb{C}}=U_{0}({\mathfrak{g}})\subset U_{1}({\mathfrak{g}})\subset U_{2}({\mathfrak{g}})\subset\cdots U_{i}({\mathfrak{g}})\subset\cdots$ (5.1)
@@
Let ${\rm gr}(U({\mathfrak{g}}))=\bigoplus_{i=0}^{\infty}U_{i}({\mathfrak{g}})/U_{i-1}({\mathfrak{g}})$ (where $U_{-1}({\mathfrak{g}}):=0$). It is called the associated graded algebra with respect to the canonical filtration. Let $\mathcal{S}({\mathfrak{g}})$ denote the supersymmetric algebra of ${\mathfrak{g}}$. We have the following superalgebra isomorphism (see [Dix, Proposition 2.3.6]):
${\rm gr}\!\left(U({\mathfrak{g}})\right)\cong\mathcal{S}({\mathfrak{g}}).$ (5.2)
@@
[…]
@@
$\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,z}\left(A\otimes t^{-1}\right)=-\sum_{i=1}^{\ell}\sum_{k=0}^{\gamma_{i}-1}\frac{A\otimes\overline{t}_{z_{i}}^{k}}{(z-z_{i})^{k+1}},\qquad\mbox{for $A\in{\mathfrak{g}}$.}$ (5.3)
@@ 5.2. Classical Gaudin algebras for ${\mathfrak{gl}}_{d}$
¯$\overline{\mathcal{L}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big{[}\delta_{i,j}w+\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,z}\left(e_{ij}\otimes t^{-1}\right)+\mu(e_{ij})\Big{]}_{i,j=1,\ldots,d}$
@@
for some $\overline{a}_{i}(z)\in\mathcal{S}\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}[\kern-1.1pt[z^{-1}]\kern-1.1pt]$. Let $\overline{\mathcal{A}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ be the subalgebra of $\mathcal{S}\big{(}{\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big{)}$ generated by the coefficients of the series $\overline{a}_{i}(z)$ for $i=1,\ldots,d$. It is called the classical Gaudin algebra for ${\mathfrak{gl}}_{d}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, or simply the classical Gaudin algebra corresponding to $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$. The algebra $\overline{\mathcal{A}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ is known to be Poisson-commutative [FFTL].
@@
${\rm gr}\!\left(\mathcal{A}_{d}^{0}({\bf z},{\boldsymbol{\gamma}})\right)\cong\overline{\mathcal{A}}_{d}^{0}({\bf z},{\boldsymbol{\gamma}}).$ (5.4)
@@
$\hskip 28.45274pt{\rm gr}(\mathcal{A}_{d}^{\mu})\cong\overline{\mathcal{A}}_{d}^{\mu}\quad\qquad\mbox{for any $\mu\in{\mathfrak{gl}}_{d}^{*}$.}$ (5.5)
@@
¯$\overline{\mathcal{L}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big{[}\delta_{i,j}z+(-1)^{|i|}\left(\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,w}\left(E^{i}_{j}\otimes t^{-1}\right)+\mu(E^{i}_{j})\right)\Big{]}_{i,j\in\mathbb{I}},$
@@
for some $\overline{b}_{i}(w)\in\mathcal{S}\big{(}{\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big{)}(\!(w^{-1})\!)$. Let $\overline{\mathcal{A}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ be the subalgebra of $S\big{(}{\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big{)}$ generated by the coefficients of the series $\overline{b}_{i}(w)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. We call ${\overline{\mathcal{A}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})}$ the classical Gaudin algebra for ${\mathfrak{gl}}_{\diamondsuit}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, or simply the classical Gaudin algebra corresponding to $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$. It is Poisson-commutative by an argument parallel to the proof of [MR, Corollary 3.6]. If $q=n=0$ and $d=p+m$, the algebra coincides with the one given in Section 5.2.
Let ${\rm gr}\!\left(\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})\right)$ be the associated graded algebra of $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ with respect the filtration induced by the canonical filtration on $U({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}}))$. We conjecture that the Gaudin algebras are quantizations of the classical Gaudin algebras.
@@
Let ${d^{\prime}}$, ${\boldsymbol{\ell}}$, ${\bf w}$, ${\bf z}$, ${\boldsymbol{\xi}}$ and ${\boldsymbol{\gamma}}$ be as in Section 4.1. In Theorem 4.5, we establish the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{\diamondsuit})$ for Gaudin models with irregular singularities, which is an equivalence between the actions of the Gaudin algebras $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$ on the Fock space ${\mathcal{F}}$. The duality and Conjecture 5.1 lead us to expect that there should be a duality between the classical Gaudin algebras corresponding to $\mathcal{A}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})$ and $\mathcal{A}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})$. In this subsection, we will see that the expectation is confirmed.
@@
$\mathcal{D}_{0}\subset\mathcal{D}_{1}\subset\mathcal{D}_{2}\subset\cdots\mathcal{D}_{k}\subset\cdots$ (5.6)
@@
${\rm gr}\!\left(\mathcal{D}\right)\cong\overline{\mathcal{D}}$ (5.7)
[…]
$\{x^{a}_{i},x^{b}_{j}\}=\{x^{a}_{i},y^{b}_{s}\}=\{y^{a}_{r},y^{b}_{s}\}=\{p_{x^{a}_{i}},p_{x^{b}_{j}}\}=\{p_{x^{a}_{i}},p_{y^{b}_{s}}\}=\{p_{y^{a}_{r}},p_{y^{b}_{s}}\}=0,$ (5.8)
$\{p_{x^{a}_{i}},x^{b}_{j}\}=\delta_{i,j}\delta_{a,b},\quad\{p_{x^{a}_{i}},y^{b}_{s}\}=\{p_{y^{a}_{r}},x^{b}_{j}\}=0,\quad\{p_{y^{a}_{r}},y^{b}_{s}\}=\delta_{r,s}\delta_{a,b},$ (5.9)
@@
¯$\overline{\mathcal{L}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\overline{\mathcal{L}}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})\quad\text{and}\quad\overline{\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=\overline{\mathcal{L}}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}}).$
@@
Recall $[\gamma_{i}]$, $1\leq i\leq{\boldsymbol{\ell}}$, defined in (4.16). We have
@@
¯$\overline{\mathcal{L}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\bigoplus_{a=1}^{{d^{\prime}}}\big{(}\!-{\rm J}_{\xi_{a}}(w_{a}-w)\big{)}-\left[\sum_{i=1}^{{\boldsymbol{\ell}}}\sum_{k=0}^{\gamma_{i}-1}\frac{e_{ab}\otimes\overline{t}_{z_{i}}^{k}}{(z-z_{i})^{k+1}}\!\right]_{\!a,b=1,\ldots,d}$
@@
¯$\overline{\mathcal{L}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=\bigoplus_{i=1}^{{\boldsymbol{\ell}}}\big{(}\!-{\rm J}_{\gamma_{i}}(z_{i}-z)\big{)}-\left[\sum_{a=1}^{{d^{\prime}}}\sum_{k=0}^{\xi_{a}-1}\frac{E^{i}_{j}\otimes\overline{t}_{w_{a}}^{k}}{(w-w_{a})^{k+1}}\!\right]_{\!i,j\in\mathbb{I}}.$
@@
Acknowledgments. The first author was partially supported by NSTC grant 113-2115-M-006-010 of Taiwan. The second author was partially supported by NSTC grant 112-2115-M-006-015-MY2 of Taiwan.

### v2

@@ 5.1. Settings
(5.1) ${\mathbb{C}}=U_{0}({\mathfrak{g}})\subset U_{1}({\mathfrak{g}})\subset U_{2}({\mathfrak{g}})\subset\cdots U_{i}({\mathfrak{g}})\subset\cdots$
@@
Let ${\rm gr}(U({\mathfrak{g}}))=\bigoplus_{i=0}^{\infty}U_{i}({\mathfrak{g}})/U_{i-1}({\mathfrak{g}})$ (where $U_{-1}({\mathfrak{g}}):=0$). It is called the associated graded algebra of $U({\mathfrak{g}})$ with respect to the canonical filtration. Let $\mathcal{S}({\mathfrak{g}})$ denote the supersymmetric algebra of $\mathfrak{g}$}. We have the following superalgebra isomorphism (see [Dix, Proposition 2.3.6]):
(5.2) ${\rm gr}\!\left(U({\mathfrak{g}})\right)\cong\mathcal{S}({\mathfrak{g}}).$
@@
A Poisson superalgebra $\mathcal{A}$ is a supercommutative superalgebra over $\mathbb{C}$} equipped with a $\mathbb{C}$}-bilinear map $\{\cdot,\cdot\}:\mathcal{A}\times\mathcal{A}\longrightarrow\mathcal{A}$, called a Poisson bracket, such that $(\mathcal{A},\{\cdot,\cdot\})$ is a Lie superalgebra satisfying the Leibniz rule, i.e.,
@@
(5.3) $\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,z}\left(A\otimes t^{-1}\right)=-\sum_{i=1}^{\ell}\sum_{k=0}^{\gamma_{i}-1}\frac{A\otimes\overline{t}_{z_{i}}^{k}}{(z-z_{i})^{k+1}},\qquad\mbox{for $A\in{\mathfrak{g}}$.}$
@@ 5.2. Classical Gaudin algebras for ${\mathfrak{gl}}_{d}$
$\overline{{\mathcal{L}}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big[\delta_{i,j}w+\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,z}\left(e_{ij}\otimes t^{-1}\right)+\mu(e_{ij})\Big]_{i,j=1,\ldots,d}$
@@
for some $\overline{a}_{i}(z)\in\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)[\kern-1.1pt[z^{-1}]\kern-1.1pt]$. Let $\overline{\mathcal{A}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ be the subalgebra of $\mathcal{S}\big({\mathfrak{gl}}_{d}({{\bf z},{\boldsymbol{\gamma}}})\big)$ generated by the coefficients of the series $\overline{a}_{i}(z)$ for $i=1,\ldots,d$. It is called the classical Gaudin algebra for ${\mathfrak{gl}}_{d}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, relative to $\mu$, or simply the classical Gaudin algebra corresponding to $\mathcal{A}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$. The algebra $\overline{\mathcal{A}}_{d}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ is known to be Poisson-commutative [FFTL].
@@
(5.4) ${\rm gr}\!\left(\mathcal{A}_{d}^{0}({\bf z},{\boldsymbol{\gamma}})\right)\cong\overline{\mathcal{A}}_{d}^{0}({\bf z},{\boldsymbol{\gamma}}).$
@@
(5.5) $\hskip 28.45274pt{\rm gr}(\mathcal{A}_{d}^{\mu})\cong\overline{\mathcal{A}}_{d}^{\mu}\quad\qquad\mbox{for any $\mu\in{\mathfrak{gl}}_{d}^{*}$.}$
@@
$\overline{{\mathcal{L}}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})=\Big[\delta_{i,j}z+(-1)^{|i|}\left(\overline{\Psi}_{({\bf z},{\boldsymbol{\gamma}}),\,w}\left(E^{i}_{j}\otimes t^{-1}\right)+\mu(E^{i}_{j})\right)\Big]_{i,j\in\mathbb{I}},$
@@
for some $\overline{b}_{i}(w)\in\mathcal{S}\big({\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big)(\!(w^{-1})\!)$. Let $\overline{\mathcal{A}}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ be the subalgebra of $S\big({\mathfrak{gl}}_{\diamondsuit}({{\bf z},{\boldsymbol{\gamma}}})\big)$ generated by the coefficients of the series $\overline{b}_{i}(w)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. It is called the classical Gaudin algebra for $\mathfrak{gl}$$_{\diamondsuit}$ with singularities of orders $\gamma_{i}$ at $z_{i}$, $i=1,\ldots,\ell$, relative to $\mu$, or simply the classical Gaudin algebra corresponding to $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$. It is Poisson-commutative by an argument parallel to the proof of [MR, Corollary 3.6]. If $q=n=0$ and $d=p+m$, the algebra coincides with the one given in Section 5.2.
Let ${\rm gr}\!\left(\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})\right)$ be the associated graded algebra of $\mathcal{A}_{\diamondsuit}^{\mu}({\bf z},{\boldsymbol{\gamma}})$ with respect to the filtration induced by the canonical filtration on $U({\mathfrak{gl}}_{\diamondsuit}({{\bf w},{\boldsymbol{\xi}}}))$. We conjecture that the Gaudin algebras are quantizations of the classical Gaudin algebras.
@@
[…]
@@
(5.6) $\mathcal{D}_{0}\subset\mathcal{D}_{1}\subset\mathcal{D}_{2}\subset\cdots\mathcal{D}_{k}\subset\cdots$
@@
(5.7) ${\rm gr}\!\left(\mathcal{D}\right)\cong\overline{\mathcal{D}}$
sending the elements $x^{a}_{i}+\mathcal{D}_{0}$, $y^{a}_{r}+\mathcal{D}_{0}$, $\partial_{x^{a}_{i}}+\mathcal{D}_{0}$ and $\partial_{y^{a}_{r}}+\mathcal{D}_{0}$ in $\mathcal{D}_{1}/\mathcal{D}_{0}$, respectively, to the elements $x^{a}_{i}$, $y^{a}_{r}$, $p_{x^{a}_{i}}$ and $p_{y^{a}_{r}}$ in $\overline{\mathcal{D}}$, for $1\leq i\leq m+n$, $1\leq r\leq p+q$ and $1\leq a\leq d$ (see, for example, [CG, Example 1.3.3] and the subsequent discussion). Moreover, the superalgebra ${\rm gr}\!\left(\mathcal{D}\right)$ is naturally a Poisson superalgebra (which is seen as in the case of ${\rm gr}\!\left(U({\mathfrak{g}})\right)$ in Section 5.1), and thus we have a Poisson superalgebra structure on $\overline{\mathcal{D}}$ via (5.7). More explicitly, $\overline{\mathcal{D}}$ is a Poisson superalgebra with the Poisson bracket $\{\cdot,\cdot\}$ given by
(5.8) $\{x^{a}_{i},x^{b}_{j}\}=\{x^{a}_{i},y^{b}_{s}\}=\{y^{a}_{r},y^{b}_{s}\}=\{p_{x^{a}_{i}},p_{x^{b}_{j}}\}=\{p_{x^{a}_{i}},p_{y^{b}_{s}}\}=\{p_{y^{a}_{r}},p_{y^{b}_{s}}\}=0,$
(5.9) $\{p_{x^{a}_{i}},x^{b}_{j}\}=\delta_{i,j}\delta_{a,b},\quad\{p_{x^{a}_{i}},y^{b}_{s}\}=\{p_{y^{a}_{r}},x^{b}_{j}\}=0,\quad\{p_{y^{a}_{r}},y^{b}_{s}\}=\delta_{r,s}\delta_{a,b},$
@@
$\overline{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\overline{{\mathcal{L}}}_{d}^{\mu_{{\boldsymbol{\xi}}}^{{\bf w}}}({\bf z},{\boldsymbol{\gamma}})\quad\text{and}\quad\overline{{\mathcal{L}}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=\overline{{\mathcal{L}}}_{\diamondsuit}^{\nu_{{\boldsymbol{\gamma}}}^{{\bf z}}}({\bf w},{\boldsymbol{\xi}}).$
@@
Recall $\gamma_{i}]$, $1\leq i\leq{\boldsymbol{\ell}}$, defined in (4.16). We have
@@
$\overline{{\mathcal{L}}}_{d}^{{\bf w},\,{\boldsymbol{\xi}}}({\bf z},{\boldsymbol{\gamma}})=\bigoplus_{a=1}^{{d^{\prime}}}\big(\!-{\rm J}_{\xi_{a}}(w_{a}-w)\big)-\left[\sum_{i=1}^{{\boldsymbol{\ell}}}\sum_{k=0}^{\gamma_{i}-1}\frac{e_{ab}\otimes\overline{t}_{z_{i}}^{k}}{(z-z_{i})^{k+1}}\!\right]_{\!a,b=1,\ldots,d}$
@@
$\overline{{\mathcal{L}}}_{\diamondsuit}^{{\bf z},{\boldsymbol{\gamma}}}({\bf w},{\boldsymbol{\xi}})=\bigoplus_{i=1}^{{\boldsymbol{\ell}}}\big(\!-{\rm J}_{\gamma_{i}}(z_{i}-z)\big)-\left[(-1)^{|i|}\sum_{a=1}^{{d^{\prime}}}\sum_{k=0}^{\xi_{a}-1}\frac{E^{i}_{j}\otimes\overline{t}_{w_{a}}^{k}}{(w-w_{a})^{k+1}}\!\right]_{\!i,j\in\mathbb{I}}.$
@@
Acknowledgments. The first author was partially supported by NSTC grant 113-2115-M-006-010. The second author was partially supported by NSTC grant 112-2115-M-006-015-MY2. The authors would like to thank the referee for valuable comments and suggestions.